Investing in treasury bills. An investment advisor uses the polynomial P(1+r)^10 to predict the value of a clients investment of P dollars
with an average anual return r. The accompanying graph shows historic average annual returns for the last 20 years for various asset classes. Use the historical average return to predict the value in 10 years of an investment of $10,000 in treasury bills.
I can't put the graph in but I can tell you about it
the numbers on the left go vertically from 0 up to 20 on the left . There are four vertical bars#1 is large company stocks it goes up to 16.7% #2 is long term corporate bonds it goes up to 10.3%, #3 US treasury bills ,goes up to 7.3%, and #4 inflation goes up to 3.4%.

Question 94:
How much more would the investment in Exercise 93 be worth in 10 years if the client invests in large company stocks rather than the U.S. Treasury bills?
This question is from textbook Elementary and Intermediate Algebra
: This is a tough one and we would greatly appreciate it if you could help us solve this one...please..
Question 93: Investing in treasury bills. An investment advisor uses the polynomial P(1+r)^10 to predict the value of a clients investment of P dollars
with an average anual return r. The accompanying graph shows historic average annual returns for the last 20 years for various asset classes. Use the historical average return to predict the value in 10 years of an investment of $10,000 in treasury bills.
I can't put the graph in but I can tell you about it
the numbers on the left go vertically from 0 up to 20 on the left . There are four vertical bars#1 is large company stocks it goes up to 16.7% #2 is long term corporate bonds it goes up to 10.3%, #3 US treasury bills ,goes up to 7.3%, and #4 inflation goes up to 3.4%.

Question 94:
How much more would the investment in Exercise 93 be worth in 10 years if the client invests in large company stocks rather than the U.S. Treasury bills?

Answer by vleith(954)

93) rate of return on T-bills from graph is 7.3% - 0.073
F = 10000(1+0.073)^10
F = 20230.06
94) rate or return for Large company is 16.7% = 0.167
F = 10000(1+0.167)^10
F = 46849.89
Difference is 46849.89 - 20230.06 = 26619.80
Compound interest matters
Lesson to learn? Save early, save at a good rate